$105 Quantitative Problem: Rosnan Industries 2018 and 2017 balance sheets and income statements are shown below. Balance SheIncome Statements: Sales Operating costs excluding depreciation EBITDA Depreciation and amortization EBIT Interest EBT TaxesWhat is the firms 2018 current ratio? Round your answer to two decimal places. If the industry average debt-to-total-assets
Answer #1

(a) Current Ratio = Current Assets / Current Liabilities

Current Assets in 2018 = $ 770

Current Liabilities in 2018 = $ 395

Current Ratio = 770/395

= 1.94

(b) Debt To Total assets ratio = Total Debt / Average Total Assets

Total Debt 2018 = Total Current liabilities + Total long term debt

= 395+450

= 845

Average Total Assets in 2018 = 2770

Debt to total assets ratio = 845/2770*100

= 30.50%

If the industry average is 30% then Roshnan's creditors have a smaller cushion than indicated by the industry average as its debt to total assets ratio = 30.50% which is higher/equal to 30%.

(c) Net Profit Margin = Earnings after interest & taxes (EAT) / Net sales

Earnings after interest & taxes of 2018 = 353 (Given)

Net sales = 2000

Net Profit Margin = 353/2000*100

= 17.65%

(d) False as Rosnan's debt to total assets ratio is equal to industry average.

(e) Price -Earning ratio = market price of share / EPS

market price per share = 25 (Given)

Earnings per share = Net profit available to equity holders/No. of common stock outstanding

= 353/100

= 3.53

P/E Ratio = 25/3.53

= 7.08

Under Du-pont Equation ROE = Earnings after taxes/Sales*Sales/Total Assets

= EAT/Total Assets

EAT in 2018 = 353

Total Assets = 2770

ROE = 353/2770

= 12.74%

Please check with your answer and let me know.